OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9

OpenStudy (anonymous):

@abb0t @perl @Jhannybean @StudyGurl14 @xoxoebnap_ @AriPotta @zepdrix @Gschibby

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@CoolGuy13

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

hello @StudyGurl14

OpenStudy (studygurl14):

4x - 5y = 9 -5y = -4x + 9 \(\large y=\frac{4}{5}x-\frac{9}{5}\)

OpenStudy (anonymous):

omg ty

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

ty tyty

OpenStudy (anonymous):

ty soooo much

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

can you help me with more @StudyGurl14

OpenStudy (studygurl14):

sure. new thread please and tag me

OpenStudy (anonymous):

ok